# SkyDiffusion: Street-to-Satellite Image Synthesis with Diffusion Models and BEV Paradigm

### v1

@@ Abstract
Street-to-satellite image synthesis focuses on generating realistic satellite images from corresponding ground street-view images while maintaining a consistent content layout, similar to looking down from the sky. The significant differences in perspectives create a substantial domain gap between the views, making this cross-view generation task particularly challenging. In this paper, we introduce SkyDiffusion, a novel cross-view generation method for synthesizing satellite images from street-view images, leveraging diffusion models and Bird’s Eye View (BEV) paradigm. First, we design a Curved-BEV method to transform street-view images to the satellite view, reformulating the challenging cross-domain image synthesis task into a conditional generation problem. Curved-BEV also includes a "Multi-to-One" mapping strategy for combining multiple street-view images within the same satellite coverage area, effectively solving the occlusion issues in dense urban scenes. Next, we design a BEV-controlled diffusion model to generate satellite images consistent with the street-view content, which also incorporates a light manipulation module to optimize the lighting condition of the synthesized image using a reference satellite. Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on both suburban (CVUSA & CVACT) and urban (VIGOR-Chicago) cross-view datasets, with an average SSIM increase of 14.5% and a FID reduction of 29.6%, achieving realistic and content-consistent satellite image generation. The code and models of this work will be released at https://opendatalab.github.io/skydiffusion/.
1 Introduction
@@
Cross-view synthesis involves transformations between disparate viewpoints, such as from street-view images to satellite images and vice versa. This process is crucial for applications in autonomous driving [2], 3D reconstruction [40, 7], virtual reality [3, 6], etc. The significant changes in street view and satellite perspectives result in drastically different appearances of objects, leading to a substantial disparity between different image domains. As shown in Fig 1 (a), in cross-view scenarios, building facades observed only in street view and rooftops observed only from satellite view are considered uncorrelated and drastically different features [24, 35, 47]. However, objects such as roads and lawns, which are below the viewpoint, can be observed in different views even if they undergo some deformation, making them cross-view shared correlated information. The key to the challenging task of cross-view generation lies in capturing cross-view relevant information and generating realistic views with consistent content layout.
@@
Figure 1: Illustration of the cross-view image synthesis task. (a) In the cross-view scenario, there are features unique to the street-view image, features exclusive to the satellite image, and shared correlated information observable from both views. (b) Comparison with existing methods using GANs (e.g., CUT [21]) or diffusion models (e.g., Instr-p2p [1]) for street-to-satellite image synthesis.
@@
In the field of cross-view synthesis, converting satellite images to street-view images has received extensive attention [30, 20, 43, 23]. Researchers have leveraged polar transformations [38, 31] or geo-transformation [20] to bridge the gap from the satellite to the street view domain, and then employ Generative Adversarial Networks (GANs) to generate realistic street views. This paper focuses on the less explored yet equally challenging task of synthesizing satellite images from street-view images. Some researchers have used semantic segmentation priors to aid in synthesizing satellite images, achieving notable results [24, 35]. However, the semantic maps are often unavailable in real scenarios. Existing street-to-satellite image generation also lacks transformations such as polar or geo-transformation to fit the gap from the street view to the satellite domain. Moreover, as shown in Fig 1(b), existing methods on street-to-satellite image synthesis, which mostly rely on GANs, often result in lower-quality satellite images with unrealistic textures compared to the ground truth (GT).
@@
Recently, diffusion models have gained attention for their exceptional image synthesis and rich 3D world knowledge derived from large-scale text and 2D image data [9, 45]. Many studies have explored text-guided fine-tuning of diffusion models to top-down perspective from other views [15, 14, 11, 27]. Additionally, novel view synthesis methods have been developed by training with multi-view images and camera poses [19, 29]. Although current novel view synthesis techniques demonstrate satisfactory performance at the object level, their performance in cross-view generation for complex scenes with significant viewpoint differences remains suboptimal. A possible solution is to treat the this task as an image translation task [16, 22, 1]. However, as shown in Fig 1(b), due to the domain gap of different views, the resulting images are often realistic but not actual, with significant discrepancies between the synthesized satellite images and the actual corresponding street view content.
In response to these challenges, this paper introduces SkyDiffusion, the first diffusion-model-based method for synthesizing satellite images from street-view images, similar to looking down from the sky. In SkyDiffusion, we design a curved-BEV method that transforms street-view images into the satellite view to achieve domain alignment between different views. Curved-BEV also includes a "Multi-to-One" BEV mapping method that can optimize BEV perception in densely occluded urban scenarios. Subsequently, we design a BEV-controlled diffusion model to generate satellite images consistent with the street-view content, incorporated with a light manipulation module to address the lack of inherent correlation in lighting conditions between the two views. SkyDiffusion outperforms existing state-of-the-art methods on both suburban (CVUSA and CVACT) and urban (VIGOR-Chicago) cross-view datasets, with an average SSIM increase of 14.5% and a FID reduction of 29.6%, successfully generating realistic and content-consistent satellite images.
2 Related Work
@@
Cross-view image synthesis. Cross-view image synthesis involves generating realistic scenes images with significant viewpoint changes, primarily divided into satellite-to-street [30, 20, 43, 23, 38] and street-to-satellite generation [35, 24, 25]. In this field, synthesizing satellite images from street views is a challenging yet important task. Tang [35] used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios. Regmi [24] attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies. Additionally, the performance of GANs has also limited the quality of street-to-satellite image synthesis in previous studies.
BEV transformation. BEV transformation is a perspective domain conversion technique similar to polar transformation [31, 38] and geo-transformation [20], widely used in retrieval [4], navigation [17] and localization tasks [34, 41, 32]. Traditional BEV transformation techniques, which rely on camera parameters [28] or depth estimation methods [37, 44], face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation. Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image [41] or feature level [32]. However, such methods are based on the ground plane assumption, where all mapped points are located below the camera, limiting the mapping of upper street scene features and causing information loss. Furthermore, to our knowledge, no studies have yet explored the use of BEV methods for the task of street-to-satellite image generation.
Novel view synthesis based on diffusion models. Diffusion models have shown significant potential in recent studies on novel view synthesis [11, 15, 36, 5, 39]. DreamBooth [27] and AerialDiffusion [14] utilize text guidance to achieve top-down image generation by fine-tuning diffusion models. However, it is challenging to fully encapsulate the complexity of real-world scenes using textual information alone. Zero-1-to-3 [19] and Zero 123++ [29] use multi-view images and camera poses to guide diffusion models in learning object structures and textures, enabling new viewpoint synthesis. However, these methods are limited to object-level tasks with strong viewpoint correlations and struggle to handle the significant viewpoint changes in cross-view generation. Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images [16, 1, 22, 45]. However, the substantial domain gap between street-view images and satellite images presents a significant challenge in generating content-consistent satellite images.
3 Method
As illustrated in Fig. 2, this paper introduces a novel method for synthesizing satellite images from corresponding street-view images. Our cross-view synthesis network, SkyDiffusion, initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings (Section 3.1). Subsequently, the BEV-Controlled diffusion model with light manipulation is employed for the controlled synthesis of satellite images (Section 3.2). Following this, we detail our network training process (Section 3.3).
@@ 3 Method
3.1 Curved BEV Transformation
One-to-One BEV. Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street view to the BEV plane[32, 41]. However, this assumption fails to map the content above the street-view image and results in noticeable distortion as points move away from the center area. Thus, we propose an improved curved BEV transformation to alleviate this problem and better capture cross-view relevant information.
@@ 3.1 Curved BEV Transformation
Furthermore, we can convert $P(x,y,z)$ to spherical coordinates $P(\theta,\varphi)$. Based on the equirectangular projection properties of street view images, $P(\theta,\varphi)$ corresponds to the panorama image index $P(u,v)$. This allows us to establish the mapping relationship between $P(x,y,z)$ and $P(u,v)$ from the following equation, where $H$ represents the camera height, and $w$, $h$ represent the width and height of the street image. More details can be found in the supplementary material.
@@
Following Eq. 1 and Eq. 2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods [28, 37, 44] (Fig. 3 (b)). This allows effective mapping of upper street view features, such as distant roads and buildings. Since the mapping relationship for each transformation is fixed, the computational overhead is minimal.
Multi-to-One BEV. Due to the limitations imposed by dense buildings on street view observations, the BEV sensing range is very limited in urban datasets like VIGOR [46]. To address this, we propose a "Multi-to-one" BEV mapping method that unifies multiple street view BEV mappings results $\textit{BEV}_{\mathrm{cam}_{k}}$ into a single satellite coordinate system based on the relationships of camera position (Fig. 3 (c)). Additionally, we use a simple semantic segmentation network to identify and discard the regions with severe BEV distortion, typically those obstructed by buildings. We select the street view mapping result closest to the capture point for overlapping areas following Eq. 3, where $\textit{BEV}\left(x,y\right)$ represents the unified BEV space corresponding to the satellite image, $x,y$ corresponds to the index position of the BEV plane, $\Delta x_{k}$ and $\Delta y_{k}$ represent the offset of the street-view image from the satellite center, and $x_{\mathrm{cam}_{i}}$ and $y_{\mathrm{cam}_{i}}$ represent the coordinates of camera position in the satellite coordinate system.
$\textit{BEV}\left(x,y\right)=\textit{BEV}_{\mathrm{cam}_{k}}\left(x-\Delta x_{k},y-\Delta y_{k}\right),k=\arg\min_{i}\sqrt{(x-x_{\mathrm{cam}_{i}})^{2}+(y-y_{\mathrm{cam}_{i}})^{2}}$ (3)
This allows us to synthesize large-scale scenes using street-view images that typically contain only limited local information.
3.2 BEV-Controlled Diffusion Model with Light Manipulation
Curved BEV controlled synthesis. Image diffusion models learn to progressively denoise images and generate samples from the training domain. The denoising process can occur in pixel space or in a latent space encoded from training data. Inspired by ControlNet [45], we enhance the control over the generated aerial views by replicating the SD module, enabling the transformed BEV images to effectively guide the synthesis process. Specifically, we first use a small encoder to capture the cross-view associative information of the input BEV image $I_{bev}$ such as road directions and building positions, to achieve BEV image feature embedding $c_{bev}$. Notably, considering the potential feature distortions in BEV images, we employ a spatial attention mechanism to enhance feature capture while suppressing distorted features.
@@ 3.2 BEV-Controlled Diffusion Model with Light Manipulation
Subsequently, by employing zero convolution layers and replicating the structures and weights of the encoder and middle block from the SD module, the extracted BEV features are injected into another SD module. Guided by the BEV image, the diffusion model considers cross-view associative information, synthesizing aerial views with consistent content.
Light Manipulation. Given the differences in perspectives and timing between satellite and street images, there is no direct correlation in lighting conditions. We employ the DNCM for modeling deterministic color mapping [12]. Initially, we use nDNCM to obtain the "image content information" $I_{c}$ from the preliminary synthesis results $I_{0}$ of the satellite images. Then, we input the lighting and color style $r_{s}$ obtained from the reference aerial view into sDNCM for lighting transfer.
$I_{s}=s\textit{DNCM}(\textit{nDNCM}(I_{0},I_{c}),r_{s}),\text{ where }\textit{DNCM}(I,T)=(I_{h\times w,3}\cdot P_{3,k})\cdot T_{k,k}\cdot Q_{k,3}$ (5)
$I_{c}$ and $r_{s}$ are the content information and color reference information extracted by the encoder $E$, as shown in Fig. 2. The DNCM module internally includes ${P}_{3,k}$ and ${Q}_{k,3}$, which represent the color embedding and recovery matrices, respectively. Here, $k$ represents the embedding dimension. The adaptive color mapping matrix ${T}_{k,k}$ is also obtained through the encoder.
3.3 Network Training
@@
Following previous image migration work [13, 12], we apply different LUT filters to the satellite images to simulate the results under various lighting conditions for training the light manipulation module. During training, we use L2 consistency loss $\mathcal{L}_{\text{con}}$ to maintain image content consistency and L1 reconstruction loss $\mathcal{L}_{\text{rec}}$ to learn color style transfer.
Implementation Details. We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model [26], with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 [10]. Final inference sampling uses 50 steps with the DDIM strategy[33]. Training is on eight NVIDIA A100 GPUs with a batch size of 128 for 100 epochs. In the curved BEV transformation, the parameter $\lambda$ is set to 15 and $H$ is set to 1.5. For light manipulation, the reference satellite image is randomly selected from normally toned images in the training set, as shown in Figure 8, and the DNCM mapping matrix parameter $k$ is set to 32 (details in Supplementary Material).
4 Experiments
@@ 4.1 Datasets
Suburban dataset (CVUSA & CVACT) are two commonly used cross-view datasets, containing centrally aligned aerial and street-view images from regions in the United States and Canberra, Australia, respectively, in a one-to-one mapping relationship. The data primarily consists of suburban scenes with relatively low-rise buildings and open views. The CVUSA dataset’s training and testing sets are randomly split, while the CVACT dataset is split by region.
@@
Urban dataset (VIGOR-Chicago) is a subset of the VIGOR dataset, containing cross-view image pairs collected from Chicago. In this dataset, the street panorama and satellite images are not centrally aligned, and multiple street-view images cover the same satellite image area, with overlapping regions between different satellite images, forming a many-to-one mapping relationship.
In CVUSA [42], the street-view image is 256 $\times$ 1024 pixels, while in CVACT [18] and VIGOR [46], it is 512 $\times$ 1024, all with north in the central column. Unlike previous methods that synthesize low-resolution satellite images (64 $\times$ 64 or 256 $\times$ 256 pixels), this experiment aims to meet practical needs by generating higher-resolution satellite images (512 $\times$ 512 pixels). The images obtained from street images after the Curved-BEV transformation are the same size as the satellite images.
4.2 Experimental Setup
We compare our approach on three datasets with several state-of-the-art cross-view synthesis methods, including GAN-based methods like X-Seq [24], CUT [21], and CDTE [38], and diffusion models for image translation like img2img-turbo [22] and Instruct pix2pix (Instr-p2p) [1]. We also compare with fine-tuned diffusion models for novel view generation, including DreamBooth [27], Aerial Diff [14], and HawkI [15]. Except for the fine-tuned diffusion models, which use a brief textual description, all other models use the same street-view input for a fair comparison. Following previous work [20, 38, 24], we employ commonly used metrics such as SSIM, PSNR and LPIPS to evaluate the content consistency of the synthesized images, and use FID [8] to assess the realism of images.
4.3 Comparison with State-of-the-Art Methods

### v2

@@ Abstract
Street-to-satellite image synthesis focuses on generating realistic satellite images from corresponding ground street-view images while maintaining a consistent content layout, similar to looking down from the sky. The significant differences in perspectives create a substantial domain gap between the views, making this cross-view generation task particularly challenging. In this paper, we introduce SkyDiffusion, a novel cross-view generation method for synthesizing satellite images from street-view images, leveraging diffusion models and Bird’s Eye View (BEV) paradigm. First, we design a Curved-BEV method to transform street-view images to the satellite view, reformulating the challenging cross-domain image synthesis task into a conditional generation problem. Curved-BEV also includes a "Multi-to-One" mapping strategy for leveraging multiple street-view images within the same satellite coverage area, effectively solving the occlusion issues in dense urban scenes. Next, we design a BEV-controlled diffusion model to generate satellite images consistent with the street-view content, which also incorporates a light manipulation module to make the lighting conditions of the synthesized satellite images more flexible. Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on both suburban (CVUSA & CVACT) and urban (VIGOR-Chicago) cross-view datasets, with an average SSIM increase of 13.96% and a FID reduction of 20.54%, achieving realistic and content-consistent satellite image generation. The code and models of this work will be released at https://opendatalab.github.io/skydiffusion/
Introduction
@@
Cross-view synthesis involves transformations between disparate viewpoints, such as from street-view images to satellite images and vice versa. This process is crucial for applications in urban planning (Li et al. 2023b), 3D reconstruction (Wang et al. 2021), virtual reality (Emmaneel et al. 2023), etc. The significant changes in street view and satellite perspectives result in drastically different appearances of objects, leading to a substantial disparity between different image domains. As shown in Fig 1 (a), in cross-view scenarios, building facades observed only in street view and rooftops observed only from satellite view are considered uncorrelated and drastically different features (Regmi and Borji 2018; Zhu et al. 2023). However, objects such as roads and lawns, which are below the viewpoint, can be observed in different views even if they undergo some deformation, making them cross-view shared correlated information. The key to the challenging task of cross-view synthesis lies in bridging domain gaps, capturing cross-view relevant information, and generating realistic views with consistent content layouts.
@@
Figure 1: Illustration of the cross-view image synthesis task. (a) In the cross-view scenario, there are features unique to the street-view image, features exclusive to the satellite image, and shared correlated information observable from both views. (b) Comparison of our street-to-satellite image synthesis method with existing cross-view synthesis methods (e.g., SelectionGAN (Tang et al. 2019)) or image translation methods based on diffusion models (e.g., Instr-p2p (Brooks, Holynski, and Efros 2023)).
@@
In the field of cross-view synthesis, converting satellite images to street-view images has received extensive attention (Shi et al. 2022; Lu et al. 2020; Qian et al. 2023). Researchers have leveraged polar transformations (Toker et al. 2021) or geo-transformation (Lu et al. 2020) to bridge the gap from the satellite to the street view domain, and then employ Generative Adversarial Networks (GANs) to generate realistic street views. This paper focuses on the less explored yet equally challenging task of synthesizing satellite images from street-view images. Some researchers have used semantic segmentation priors to aid in synthesizing satellite images, achieving notable results (Tang et al. 2019). However, the semantic maps are often unavailable in real scenarios. Existing street-to-satellite image generation also lacks transformations such as polar or geo-transformation to fit the gap from the street view to the satellite domain. Moreover, as shown in Fig 1(b), existing methods on street-to-satellite image synthesis, which mostly rely on GANs, often result in lower-quality satellite images with unrealistic textures.
@@
Recently, diffusion models have gained attention for their exceptional image synthesis (Ho, Jain, and Abbeel 2020; Zhang, Rao, and Agrawala 2023). Many studies have explored text-guided fine-tuning of diffusion models to top-down perspective from other views (Kothandaraman et al. 2023a; Ruiz et al. 2023). Additionally, novel view synthesis methods have been developed by training with multi-view images and camera poses (Liu et al. 2023; Shi et al. 2023a). Although current novel view synthesis techniques demonstrate satisfactory performance at the object level, their performance in cross-view generation for complex scenes with significant viewpoint differences remains suboptimal. A possible solution is to treat the this task as an image translation task (Li et al. 2023a; Parmar et al. 2024). However, as shown in Fig 1(b), due to the domain gap of different views, the resulting images are often realistic but not actual, with significant discrepancies between the synthesized satellite images and the actual corresponding street view content. In response to these challenges, this paper introduces SkyDiffusion, a street-to-satellite image synthesis method based on diffusion models and the BEV paradigm, similar to viewing from the sky. Unlike previous cross-view synthesis methods, our approach is based on a diffusion framework, enabling the generation of more realistic satellite images. We design a Curved-BEV method that transforms street-view images into satellite views, achieving domain alignment between different perspectives and facilitating content-consistent satellite image synthesis. To address the limited perception range of BEV in densely occluded urban scenarios, the Curved-BEV method also includes a Multi-to-One mapping strategy. Additionally, our BEV-controlled diffusion model incorporates a light manipulation module, making the lighting conditions of the synthesized satellite images more adaptable. Our main contributions are summarized as follows:
  • We introduce SkyDiffusion, a novel street-to-satellite image synthesis method leveraging diffusion models and BEV paradigm to generate realistic and content-consistent satellite images.
  • We design a Curved-BEV method to transform street-view images into satellite views for domain alignment. It also includes "Multi-to-One" mapping strategy to enhance BEV perception range in densely occluded urban areas.
  • SkyDiffusion outperforms state-of-the-art methods on suburban (CVUSA, CVACT) and urban (VIGOR-Chicago) cross-view datasets, with a 13.96% SSIM increase and 20.54% FID reduction.
Related Work
@@
Cross-view image synthesis involves generating realistic scenes images with significant viewpoint changes, primarily divided into satellite-to-street (Shi et al. 2022; Lu et al. 2020; Qian et al. 2023) and street-to-satellite generation (Regmi and Borji 2018, 2019). In this field, synthesizing satellite images from street views is a challenging yet important task. Regmi and Borji (2018) attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies. Tang et al. (2019) used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios. Additionally, the performance of GANs has also limited the quality of street-to-satellite image synthesis in previous studies.
@@
BEV transformation is a perspective domain conversion technique similar to polar transformation (Shi et al. 2019; Toker et al. 2021) and geo-transformation (Lu et al. 2020), widely used in retrieval (Ye et al. 2024b), navigation (Li et al. 2022) and localization tasks (Song et al. 2023; Wang et al. 2023; Shi et al. 2023b). Traditional BEV transformation techniques, which rely on camera parameters (Sarlin et al. 2023) or depth estimation methods (Teng et al. 2024; Ye et al. 2024a), face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation. Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image (Wang et al. 2023) or feature level (Shi et al. 2023b). However, such methods are based on the ground plane assumption, where all mapped points are located below the camera, limiting the mapping of upper street scene features and causing information loss. Furthermore, to our knowledge, no studies have yet explored the use of BEV methods for street-to-satellite image synthesis task.
Novel view synthesis based on diffusion models. Diffusion models have shown significant potential in recent studies on novel view synthesis (Kawar et al. 2023; Gao et al. 2024; Tseng et al. 2023). DreamBooth (Ruiz et al. 2023) and AerialDiffusion (Kothandaraman et al. 2023a) utilize text guidance to achieve top-down image generation by fine-tuning diffusion models. However, it is challenging to fully encapsulate the complexity of real-world scenes using textual information alone. Zero-1-to-3 (Liu et al. 2023) and Zero 123++ (Shi et al. 2023a) use multi-view images and camera poses to guide diffusion models in learning object structures and textures, enabling new viewpoint synthesis. However, these methods are limited to object-level tasks with strong viewpoint correlations and struggle to handle the significant viewpoint changes in cross-view generation. Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images (Li et al. 2023a; Brooks, Holynski, and Efros 2023; Zhang, Rao, and Agrawala 2023). However, the substantial domain gap between street-view images and satellite images presents a significant challenge in generating content-consistent satellite images.
Methods
As illustrated in Fig. 2, this paper introduces SkyDiffusion, a novel method for synthesizing satellite images from corresponding street-view images. It initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings. Subsequently, the BEV-Controlled diffusion model with light manipulation is employed for the controlled synthesis of satellite images. Following this, we detail our network training process.
Curved BEV Transformation
One-to-One BEV. Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street view to the BEV plane(Shi et al. 2023b; Wang et al. 2023). However, this assumption fails to map the content above the street-view image and results in noticeable distortion as points move away from the center area. Thus, we propose an enhanced curved BEV transformation to better capture cross-view information and address this issue.
@@ Methods
Furthermore, we can convert $P(x,y,z)$ to spherical coordinates $P(\theta,\varphi)$. Based on the equirectangular projection properties of street view images, $P(\theta,\varphi)$ corresponds to the panorama image index $P(u,v)$. This allows us to establish the mapping relationship between $P(x,y,z)$ and $P(u,v)$ from the following equation, where $H$ represents the camera height, and $w$, $h$ represent the width and height of the street image. More details are in the supplementary material.
@@
Following Eq. 1 and Eq. 2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods (Teng et al. 2024; Ye et al. 2024a) (Fig. 3 (b)). This effectively maps upper street view features like distant roads and buildings. Since the mapping relationship for each transformation is fixed, the computational overhead is minimal.
Multi-to-One BEV. Due to the limitations imposed by dense buildings on street view observations, the BEV sensing range is very limited in urban datasets like VIGOR. To address this, we propose a "Multi-to-one" BEV mapping method that unifies multiple street view BEV mappings results $\textit{BEV}_{\mathrm{cam}_{k}}$ into a single satellite coordinate system based on the relationships of camera position (Fig. 3 (c)). We use a simple semantic segmentation network to identify and discard the regions with severe BEV distortion, typically those obstructed by buildings. We select the street view mapping result closest to the capture point for overlapping areas following Eq. 3, where $\textit{BEV}\left(x,y\right)$ represents the unified BEV space corresponding to the satellite image, $x,y$ corresponds to the index position of the BEV plane, $\Delta x_{k}$ and $\Delta y_{k}$ represent the offset of the street-view image from the satellite center, and $x_{\mathrm{cam}_{i}}$ and $y_{\mathrm{cam}_{i}}$ represent the coordinates of camera position in the satellite coordinate system.
$&\textit{BEV}\left(x,y\right)=\textit{BEV}_{\mathrm{cam}_{k}}\left(x-\Delta x_{k},y-\Delta y_{k}\right)\\ &k=\arg\min_{i}\sqrt{(x-x_{\mathrm{cam}_{i}})^{2}+(y-y_{\mathrm{cam}_{i}})^{2}}\end{aligned}\right.$ (3)
This enables us to synthesize large-scale scenes using street-view images with limited local information.
BEV-Controlled Synthesis with Light Manipulation
Curved BEV controlled synthesis. Image diffusion models progressively denoise images and generate samples from the training domain. The denoising process can occur in pixel space or in a latent space encoded from training data. Inspired by ControlNet (Zhang, Rao, and Agrawala 2023), we enhance the control over the generated aerial views by replicating the SD module, enabling the transformed BEV images to effectively guide the synthesis process. Specifically, we first use a small encoder to capture the cross-view associative information of the input BEV image $I_{bev}$ such as road directions and building positions, to achieve BEV image feature embedding $c_{bev}$. Notably, considering the potential feature distortions in BEV images, we employ a spatial attention mechanism to enhance feature capture while suppressing distorted features.
@@
Subsequently, by using zero convolution layers and replicating the encoder and middle block structures and weights from the SD module, the extracted BEV features are injected into another SD module. Guided by the BEV image, the diffusion model incorporates cross-view associative information, synthesizing aerial views with consistent content.
Light Manipulation. Given the differences in perspectives and timing between satellite and street images, there is no direct correlation in lighting conditions. To address this, we designed a light manipulation module to make the lighting conditions of the synthesized satellite images more adaptable. We employ the DNCM for modeling deterministic color mapping (Ke et al. 2023). Initially, we use nDNCM to obtain the "image content information" $I_{c}$ from the preliminary synthesis results $I_{0}$ of the satellite images. Then, we input the lighting and color style $r_{s}$ obtained from the reference aerial view into sDNCM for lighting transfer.
{$&I_{s}=s\textit{DNCM}(\textit{nDNCM}(I_{0},I_{c}),r_{s})\\ &\textit{DNCM}(I,T)=(I_{h\times w,3}\cdot P_{3,k})\cdot T_{k,k}\cdot Q_{k,3}\end{aligned}\right.$ (5)
Figure 4: Qualitative Comparison of Synthesis Results. The first and second rows present the results on the CVUSA dataset, while the third and fourth rows display the results on the CVACT dataset. The fifth and sixth rows showcase the visualized outcomes on the VIGOR-Chicago dataset.
$I_{c}$ and $r_{s}$ are the content information and color reference information extracted by the encoder $E$, as shown in Fig. 2. The DNCM module internally includes ${P}_{3,k}$ and ${Q}_{k,3}$, which represent the color embedding and recovery matrices, respectively. Here, $k$ represents the embedding dimension. The adaptive color mapping matrix ${T}_{k,k}$ is also obtained through the encoder. Notably, the lighting adjustment module is independent of the satellite image synthesis, meaning that satellite images can be generated without relying on this module. We obtain lighting information from randomly selected satellite images in the training set, and standard images can also be used as lighting control conditions.
Network Training
@@
Following previous work (Ke et al. 2022, 2023), we apply different LUT filters to the satellite images to simulate the results under various lighting conditions for training the light manipulation module. During training, we use L2 consistency loss $\mathcal{L}_{\text{con}}$ to maintain image content consistency and L1 reconstruction loss $\mathcal{L}_{\text{rec}}$ to learn color style transfer.
Method CVUSA CVACT VIGOR-Chicago
FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
X-Seq 161.16 0.084 11.97 0.706 190.12 0.042 12.41 0.661 - - - -
CDTE 122.84 0.143 10.01 0.694 160.81 0.091 12.59 0.663 - - - -
SelectionGAN 116.57 0.129 12.38 0.742 100.21 0.116 12.04 0.684 149.53 0.127 11.77 0.778
Aerial Diff 136.18 0.103 10.06 0.855 127.29 0.108 10.24 0.878 123.16 0.141 11.49 0.831
HawkI 90.98 0.114 11.09 0.783 135.75 0.112 10.97 0.764 123.47 0.124 9.67 0.752
CUT 72.83 0.121 12.09 0.687 62.22 0.102 13.11 0.664 69.42 0.169 13.66 0.665
img2img-turbo 77.95 0.127 12.64 0.685 73.24 0.103 12.19 0.679 80.10 0.135 11.28 0.690
Instr-p2p 38.01 0.138 12.29 0.697 49.62 0.116 12.51 0.682 50.30 0.163 10.62 0.689
Ours 29.18 0.168 14.58 0.635 35.51 0.123 13.58 0.657 45.29 0.186 11.69 0.661
Table 1: Quantitative comparison of different methods on CVUSA, CVACT and VIGOR-Chicago.
Implementation Details. We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model (Rombach et al. 2022), with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 (Ho and Salimans 2022). Final inference sampling uses 50 steps with the DDIM strategy (Song, Meng, and Ermon 2021). Training is on eight NVIDIA A100 GPUs with a batch size of 128 for 100 epochs. For light manipulation, the reference satellite image is randomly selected from normally toned images in the training set, and the DNCM mapping matrix parameter $k$ is set to 32 (details in Supplementary Material).
Experiments
@@
CVUSA & CVACT (Suburban dataset). CVUSA (Workman, Souvenir, and Jacobs 2015) and CVACT (Liu and Li 2019) are two widely used cross-view datasets. They contain centrally aligned satellite and street-view images from regions in the United States and Canberra, Australia, respectively, with a one-to-one mapping relationship. The data primarily consists of suburban scenes characterized by relatively low-rise buildings and open views. The CVUSA dataset’s training and testing sets are randomly split, while the CVACT dataset is split by region.
@@
VIGOR-Chicago (Urban dataset). VIGOR-Chicago (Zhu, Yang, and Chen 2021) is a subset of the VIGOR dataset, containing cross-view image pairs collected from Chicago urban area. In this dataset, the street panorama and satellite images are not centrally aligned, and multiple street-view images cover the same satellite image area, with overlapping regions between different satellite images, forming a multi-to-one mapping relationship. All input street images are set to 512 × 1024 pixels, with the center column aligned to the north. Unlike previous methods that synthesize low-resolution satellite images (64 $\times$ 64 or 256 $\times$ 256 pixels), this experiment aims to meet practical needs by generating higher-resolution satellite images (512 $\times$ 512 pixels). In the Curved-BEV transformation, $\lambda$ is set to 3 and $H$ to 1.5 for CVUSA and CVACT datasets. For the VIGOR-Chicago dataset, $\lambda$ is set to 10, with $H$ unchanged. The Curved-BEV transformation results in street images that match the size of the satellite images.
Experimental Setup
We compared our method on three datasets with several state-of-the-art cross-view synthesis methods, including the commonly used cross-view satellite synthesis methods X-Seq (Regmi and Borji 2018), SelectionGAN (without semantic masks) (Tang et al. 2019), and CDTE (Toker et al. 2021), image translation methods such as the GAN-based CUT (Park et al. 2020) and diffusion-based Instruct pix2pix (Instr-p2p) (Brooks, Holynski, and Efros 2023) and img2img-turbo (Parmar et al. 2024), as well as novel view synthesis methods like Aerial Diff (Kothandaraman et al. 2023a) and HawkI (Kothandaraman et al. 2023b).
Except for the fine-tuned diffusion model, which requires brief textual descriptions, all other models use the same street-view input for a fair comparison.
Following previous work (Lu et al. 2020; Toker et al. 2021; Regmi and Borji 2018), we employ commonly used metrics such as SSIM, PSNR and LPIPS to evaluate the content consistency of the synthesized images, and use FID (Heusel et al. 2017) to assess the realism of images.
Comparison with State-of-the-Art Methods